Consider a buffer solution made up of H2PO4- and HPO42-. What mole ratio of HPO42- / H2PO4- will give a pH of 6.95? Ka = 6.3 x 10-8

1
Expert's answer
2021-09-09T01:52:02-0400

pH = pKa + Log10(base/acid)

-log 6.3 × 10-8 = 7.3

Therefore, log10(base/acid) = pH-pKa

= 6.95 -7.3 = -0.08

= 100.08 = 0.832

or 1/0.832 = 1.2